Part of the proceeds from a garage sale was 445.00 dollars worth of $5 and $20 bills. If there were 4 more $5 bills than $20 bills, find the number of each denomination.
Answer by sofiyac(695)

lets say we have x $20 bills
then we have x+4 $5 bills
20x+5(x+4)=445
20x+5x+20=445
25x=425
x=17
so we have 17 $20 bills
and 17+4=21 $5 bills